OpenStudy (anonymous):

The half-life of carbon- 14 is 5600 years. If charred logs from an old log cabin sow only 79% of the carbon-14 expected in living matter, when did the cabin burn down? Assume that the cabin burned soon after it was built from freshly cut logs. How many years ago did the cabin burn down?

OpenStudy (anonymous):

I got 1904 years

OpenStudy (tkhunny):

Okay, let's see how you achieved that result.

OpenStudy (anonymous):

ln(.79)/ln(20)*-5600=1904

OpenStudy (tkhunny):

I'm just not sure what that means. We have \(e^{-k(5600)} = 0.5\). This leads to \(-5600k = ln(0.5) = -ln(2)\) or \(k = \dfrac{ln(2)}{5600}\) Did you do anything like that?